OpenStudy (anonymous):

f(x)= 3x^2-3x Rewrite the quadratic function in intercept form--show steps?

OpenStudy (anonymous):

\[f(x) = 3(x^{2}-1x+1/4) -3/4\]

OpenStudy (anonymous):

\[f(x)=3(x-\frac{1}{2})^{2}-3/4\]